OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (anonymous):

y - y1 = m (x-x1)

OpenStudy (anonymous):

you can use the formula (y-y1) = m (x-x1) where m=-3 and x1,y1= (2,-5)